OpenStudy (anonymous):

Please HELP What is the value of x? 3(x-2)+5x=18 x=?

OpenStudy (like1234):

3(x-2)+5x=18 Simple and best practice solution for 3(x-2)+5x=18 equation. Check how easy it is, and learn it for the future. Our solution is simple, and easy to understand, so don`t hesitate to use it as a solution of your homework. If it's not what You are looking for type in the equation solver your own equation and let us solve it. Equation: Solution for 3(x-2)+5x=18 equation: Simplifying 3(x + -2) + 5x = 18 Reorder the terms: 3(-2 + x) + 5x = 18 (-2 * 3 + x * 3) + 5x = 18 (-6 + 3x) + 5x = 18 Combine like terms: 3x + 5x = 8x -6 + 8x = 18 Solving -6 + 8x = 18 Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '6' to each side of the equation. -6 + 6 + 8x = 18 + 6 Combine like terms: -6 + 6 = 0 0 + 8x = 18 + 6 8x = 18 + 6 Combine like terms: 18 + 6 = 24 8x = 24 Divide each side by '8'. x = 3 Simplifying x = 3

OpenStudy (like1234):

http://www.geteasysolution.com/3(x-2)+5x=18

OpenStudy (anonymous):

3x-6+15x=18 18x-6=18 +6 +6 18x=24 divide by 18 on each side x=1.3